OpenStudy (starryknight):

please explain in steps Find the sum of the first 28 terms of the following sequence: -7, -4, -1 . . . 938 1876 134 588

OpenStudy (solomonzelman):

a1=-7 d=3 \(\large\color{blue}{ \rm a_{n}=a_1+d(n-1)}\) \(\large\color{blue}{ \rm a_{28}=-7+3(28-1)}\)

OpenStudy (solomonzelman):

after you find the a28, \(\large\color{blue}{ \rm (a_1+a_{28}) \times (1/2) \times 28}\)

OpenStudy (starryknight):

i tried doing it and got 1330 using s(n)=(2a+(n-1)d)

OpenStudy (starryknight):

the guy that helped me told me to use that

OpenStudy (starryknight):

28/2(14+(28-1)(-3)

OpenStudy (starryknight):

is that how you do it

OpenStudy (solomonzelman):

I get 74 for a28. and S28=1036

OpenStudy (solomonzelman):

fisrt you have to find a28, *as I told you) and then plug in (a1 and a28) into the sum formula.